A and X are the matrices
a b
c d !
and
x y
u v !
respectively, where b is not equal
to zero. Prove that if AX = XA then u = cy/b and v = x + (d − a)y/b. Hence prove
that if AX = XA then there are numbers p and q such that X = pA + qI, and find
p and q in terms of a, b, x, y.
1
Expert's answer
2020-05-31T15:10:45-0400
"A=\begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix},X=\begin{pmatrix}
 x & y\\
 u & v
\end{pmatrix}"

"AX=\begin{pmatrix}
 a & b\\
 c & d
\end{pmatrix}\begin{pmatrix}
 x & y\\
 u & v
\end{pmatrix}=\begin{pmatrix}
 ax+bu & ay+bv\\
 cx+du & cy+dv
\end{pmatrix}"

"XA=\begin{pmatrix}
 x & y\\
 u & v
\end{pmatrix}\begin{pmatrix}
 a & b\\
 c & d
\end{pmatrix}=\begin{pmatrix}
 ax+cy & bx+dy\\
 au+cv & bu+dv
\end{pmatrix}"

"AX=XA=>"

"\begin{pmatrix}
 ax+bu & ay+bv\\
 cx+du & cy+dv
\end{pmatrix}=\begin{pmatrix}
 ax+cy & bx+dy\\
 au+cv & bu+dv
\end{pmatrix}"

"\begin{matrix}
 ax+bu=ax+cy \\
 ay+bv=bx+dy \\
 cx+du=au+cv \\
 cy+dv=bu+dv
\end{matrix}"

Then


"bu=cy=>u={cy\over b},b\not=0"

"bv=bx+(d-a)y=>v=x+{d-a\over b}y,b\not=0"

"X=pA+qI"

"X=\begin{pmatrix}
 x & y\\
 u & v
\end{pmatrix}=\begin{pmatrix}
 x & y\\
 {cy \over b} &x+ {d-a \over b}y
\end{pmatrix}="

"=p\begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}+q\begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix}"

"\begin{matrix}
 pa+q=x \\
 pb=y \\
 pc={cy \over b} \\
 pd+q=x+{d-a \over b}y
\end{matrix}"

"p={y \over b}"

"q=x-{a \over b}y"